A pot contains 2 L of brine at a concentration of 90 g/L. How much of the water should be boiled off to increase the concentration to 180 g/L
Answer by ptaylor(2198)

Let x=amount of water that needs to be boiled off
Now we know that the total amount of salt that exists before before x amount of water is boiled off has to equal the total amount of salt that exists after the water is boiled off.
Total amount of salt before the water is boiled off is (2 liters)*90 g/l=180g
Total amount of salt after x amount of water is boiled off is (2-x)liters*180g/l
Soooo
180g=(2-x)180g/l
180=360-180x
-180x=-180
x=1 liter
Hope this helps ---ptaylor